OpenStudy (anonymous):

Find the product of (x − 3)^2 Find the product of (x + 3)^2.

OpenStudy (blurbendy):

(x-3)(x-3) = x^2 -6x +9 (x + 3)(x + 3) = x^2 +6x + 9

OpenStudy (anonymous):

thank you!!

OpenStudy (blurbendy):

you are welcome!